Before arc welding was developed, a displacement reaction involving aluminum and iron(III) oxide was commonly used to produce molten iron (the thermite process). This reaction was used, for example, to connect sections of iron railroad track.


Calculate the mass of molten iron produced when 2.30 kg of aluminum reacts with 19.6 mol of iron (III) oxide.


Answer in scientific notation.


1
Expert's answer
2021-02-12T05:00:17-0500

Fe2O3 + 2Al "\to" 2Fe + Al2O3

moles of Aluminum = 2300/27 = 85.18 moles

So Fe2O3 will act as limiting reactant

and hence 19.6 moles of Fe2O3 react with Aluminum to produce 39.2 moles

mass of iron produce = 39.2×56 = 2195.2g